# The Honeycomb Conjecture in normed planes and an alpha-convex variant of a theorem of Dowker

### v1

@@ 4. Results about not necessarily polygonal norms
In this section we collect our results about the $\alpha$-honeycomb properties of general norms, and the $\alpha$-Dowker properties of general convex disks. Our first result shows sufficiently smooth convex disks ‘asymptotically’ satisfy the log-Dowker property.
@@
$f(n)=\frac{2AB(n^{4}+1)}{(n^{2}-1)^{2}n^{2}}+o\left(\frac{1}{n^{4}}\right),$
@@
Recall that $A_{\mathbf{B}^{2}}(n)=n\tan\frac{\pi}{n}$. In the following theorem, we let
Theorem 4.3.
Proof.
$\frac{1}{1+\varepsilon}\left(\frac{n-6}{n-m}\sqrt{m\tan\frac{\pi}{m}}+\frac{6-m}{n-m}\sqrt{n\tan\frac{\pi}{n}}\right)\geq\frac{2}{1-\varepsilon}\sqrt{6\tan\frac{\pi}{6}}.$
in Section 3 we have seen that there are polygons which do not satisfy the $\alpha$-Dowker property for any $\alpha<1$. Our next theorem can be regarded as a counterpoint of this observation.
@@
First, observe that by our conditions, any supporting line of $K$ intersects $K$ in a unique point, and for any point $p$ of $\operatorname{bd}(K)$ there is a unique supporting line of $K$ that contains $p$. This yields, in particular, that for any $n\geq 3$ and any minimum area convex $n$-gon $Q_{n}$ circumscribed about $K$, $Q_{n}$ has exactly $n$ sides, and the tangent points on those sides are distinct. For any point $p\in\operatorname{bd}(K)$, we denote the unique closed half plane supporting $K$ at $p$ by $H(p)$, and the boundary of $H(p)$ by $L(p)$. We call boundary points of $K$ with opposite unit normal vectors opposite, or antipodal points. Note that if $p,q\in\operatorname{bd}(K)$ are not antipodal points of $K$, there there is a unique closed arc of $\operatorname{bd}(K)$ connecting them that does not contain antipodal points. We denote this arc by $\widehat{pq}$, and denote by $A(\widehat{pq})$ the area of the bounded, connected region $R(\widehat{pq})$ whose boundary consists of $\widehat{pq}$ and the segments of $L(p)$ and $L(q)$ connecting $p,q$ to the intersection point of the two lines. We observe that the problem of finding $A_{K}(n)$ coincides with the problem of finding a tiling of $\operatorname{bd}(K)$ by $n$ arcs $\widehat{x_{i}x_{i+1}}$, $i=1,2,n$, $x_{n+1}=x_{i}$ such that $\sum_{i=1}^{n}A(\widehat{x_{i},x_{i+1}})$ is minimal.
First, we fix an arbitrary $n\geq 4$, and show that $A_{K}(n-1)+A_{K}(n+1)>2A_{K}(n)$. To do it, we follow the proof of the original theorem of Dowker [8, Theorem 1]
@@
Observe that the family of arcs $\widehat{p_{i}p_{i+1}}$, $\widehat{q_{j}q_{j+1}}$ is a $2$-tiling $\mathcal{T}_{0}$ of $\operatorname{bd}(K)$; that is every point of $\operatorname{bd}(K)$ belongs to at least two such arcs, and no point belongs to the interior of more than $2$ such arcs. Our main goal will be to modify the $2$-tiling $\mathcal{T}_{0}$ in such a way that the value of $f$ does not decrease but the number of covering pairs strictly decreases.
Note that since $\mathcal{T}_{0}$ is the union of two tilings consisting of $(n-1)$ and $(n+1)$ arcs, respectively, $\mathcal{T}_{0}$ contains covering pairs. Assume that $\widehat{v_{i}v_{j}}$ covers $\widehat{v_{k}v_{l}}$. Then let $\mathcal{T}_{1}$ denote the $2$-tiling of §$\S^{1}$ in which $\widehat{v_{i}v_{j}}$ and $\widehat{v_{k}v_{l}}$ are replaced by $\widehat{v_{i}v_{l}}$ and $\widehat{v_{k}v_{j}}$. Our main observation, as in the proof of [8, Theorem 1] is that in this case
@@
According to our conditions, $\sum_{S\in\mathcal{T}_{0}}A(S)\leq\sum_{S\in\mathcal{T}_{1}}A(S)$. Furthermore, as $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ are $2$-tilings, and $\widehat{v_{k}v_{l}}$ is already covered twice by the arcs considered in the modification, the number of covering pairs in $\mathcal{T}_{1}$ is strictly less than in $\mathcal{T}_{0}$. Repeating this procedure we obtain a $2$-tiling $\mathcal{T}_{m}$ of $\operatorname{bd}(K)$ for which $\sum_{S\in\mathcal{T}_{0}}A(S)\geq\sum_{S\in\mathcal{T}_{,}}A(S)$ and which does not contain covering pairs. Then, $A_{t}$ decomposes into two tilings $\{\widehat{v_{1},v_{3}},\widehat{v_{3}v_{5}},\ldots,\widehat{v_{2n-1}v_{1}}\}$ and $\{\widehat{v_{2},v_{4}},\widehat{v_{4}v_{6}},\ldots,\widehat{v_{2n}v_{2}}\}$, each of which contains exactly $n$ arcs.
@@
Since $V_{1}$ and $V_{2}$ are convex $n$-gons circumscribed about $K$, if in (4.2) we have strict inequality, or if $\operatorname{area}(V_{i})>A_{K}(n)$ for $i=1$ or $i=2$, we are done. Thus, in the following we assume for contradiction that $\operatorname{area}(Q_{n-1})+\operatorname{area}(Q_{n+1})=\operatorname{area}(V_{1})+\operatorname{area}(V_{2})$, and $\operatorname{area}(V_{1})=\operatorname{area}(V_{2})=A_{K}(n)$. Then, by the remark right after (4.1) about the equality case in the inequality, there is no arc $\widehat{p_{i}p_{i+1}}$ that contains points $q_{j},q_{j+1}$ in its interior, and vice versa. In particular, there are at least two vertices of $P_{n-1}$ that coincide with some vertices of $P_{n+1}$.
@@
Now, let us call a point $q_{j}$ (resp. $p_{i}$) a double vertex, if it coincides with some $p_{i}$ (resp. $q_{j}$). If for some $j$, both $q_{j}$ and $q_{j+1}$ are double vertices, than $q_{j}$ is a double vertex for every $j$ by the previous consideration, which contradicts the fact that $Q_{n+1}$ has strictly more vertices than $Q_{n-1}$. If there is no value of $j$ such that both $q_{j}$ and $q_{j+1}$ are double vertices, then by the previous consideration any arc of $\operatorname{bd}(K)$ connecting two consecutive double vertices contain the same number of vertices from $Q_{n-1}$ and $Q_{n+1}$. But this also contradicts the facet that $Q_{n+1}$ has strictly more vertices than $Q_{n-1}$.
@@ 5. A conjecture of Steinhaus
Furthermore, if $\mathcal{M}$ is a Euclidean plane, that $\mathcal{T}_{hex}$ is a regular hexagonal tiling.
@@
The first part of Theorem 5.2 follows from a straightforward modification of the proof of Theorem 1.9. For the second statement we remark that, according to his proof, the cells of $\mathcal{T}_{hex}$ are homothetic copies of a minimum area hexagon circumscribed about the isoperimetrix of the norm. Thus, if the norm is Euclidean, then these hexagons are regular. ∎
@@
For any normed plane $\mathcal{M}$ there is a hexagonal tiling $\mathcal{T}_{hex}$ of $\mathcal{M}$ such that for any convex tiling $\mathcal{T}$ of $\mathcal{M}$ satisfying $\overline{v}(\mathcal{T})\leq 6$ (resp. $\underline{v}(\mathcal{T})\leq 6$), we have
$\underline{I}(\mathcal{T})\geq I(\mathcal{T}_{hex})\quad(\hbox{resp. }\overline{I}(\mathcal{T})\geq I(\mathcal{T}_{hex})).$
@@
Question 5.5.
Let $\mathcal{T}$ be a convex tiling of $\mathbb{R}^{2}$ such that there are universal constants $\lambda,\mu>0$ such that for every cell $C$ of $\mathcal{T}$, we have $\operatorname{ir}(C)\geq\lambda$ and $\frac{\operatorname{cr}(C)}{\operatorname{ir}(C)}\leq\mu$. Is it true that $\overline{v}(\mathcal{T})\leq 6$ (resp. $\underline{v}(\mathcal{T}))\leq 6$?
@@
We define a tiling $\mathcal{T}$ of $\mathbb{R}^{2}$ in consecutive steps. More specifically, we start with the family $\mathcal{T}_{0}=\{\mathbf{S}\}$, and for every positive integer $k$, in the $k$th step step we add some cells to the family $\mathcal{F}_{k-1}$ of the already defined cells to tet the family $\mathcal{F}_{k}$. We will do it in such a way that after each step, $\bigcup\mathcal{F}_{k}=N_{k}\mathbf{S}$ for some positive integer $N$. Each time we use one of Step A and Step B, described below:
Step A: Let $\bigcup\mathcal{F}_{k-1}=N_{k-1}\mathbf{S}$, and let the sides of this square be $E_{1},E_{2},E_{3},E_{4}$. Consider a rectangle $R$ of side lengths $N_{k-1}$ and $2N_{k-1}$. We set $\mathcal{T}_{k}=\mathcal{T}_{k-1}\cup\{R_{1},R_{2},R_{3},R_{4}\}$, where each $R_{i}$ is congruent to $R$, and one half of a longer side of $R_{i}$ is $E_{i}$. We do it in such a way that the rectangles and $\bigcup\mathcal{T}_{k-1}$ are pairwise nonoverlapping (see Figure 3). Step B: Let $\bigcup\mathcal{F}_{k-1}=N_{k-1}\mathbf{S}$. We fill the region obtained as the closure of $(N_{k-1}+2)\mathbf{S}\setminus(N_{k-1}\mathbf{S})$ by $44N_{k-1}+$ 4 translates of $\mathbf{S}$.
@@
Fix any positive number $\nu>0$. First, consider the case that we apply Step A $k_{1}$ times, and after that we apply Step B once. Then $N_{k_{1}}=3^{k_{1}}$, and thus, $\mathcal{T}_{k_{1}}$ has $43^{k_{1}}$ boundary edges. In addition, $\mathcal{T}_{k_{1}}$ consists of $14k_{1}+$ 1 cells. We count the number of interior edges: Note that $\mathcal{T}_{0}$ has $4$ edges. Then in each step before the $k_{1}$th one we add $12$ edges, which are all interior edges of $\mathcal{T}_{k_{1}}$. Finally, in the $k_{1}$th step we add four interior edges of $\mathcal{T}_{k_{1}}$. Thus, in total, $\mathcal{T}_{k_{1}}$ has $412k_{1}-$ 4 interior edges. Now, if we set $R_{1}=N_{k_{1}}+\frac{1}{2}$, then
$v_{\square}(R_{1})=\frac{\sum_{C\in\mathcal{T}_{\square}(R_{1})}v(C)}{\operatorname{card}(\mathcal{T}_{\square}(R_{1}))}=\frac{43^{k_{1}}+2(12k_{1}-4)}{4k_{1}+1}.$
@@ 6. An additional remark
Estimates regarding the numbers of sides of tilings has been a longstanding problem of geometry. The classical book of Grünbaum and Shephard [14] shows that it is possible to tile the Euclidean plane with convex hexagons, but that any normal tiling contains infinitely many cells with at most six sides. Here we recall that a (not necessarily convex) tiling $\mathcal{T}$ of $\mathbb{R}^{2}$ is called normal if
@@
    the cells are universally bounded, i.e. there are universal constants $0<\hat{r}<\hat{R}$ such that every cell contains a ball of radius $\hat{r}$ and is contained in a ball of radius $\hat{R}$.
The statement of Grünbaum and Shephard was reproved in a simpler way by Kazanci and Vince in [16]. For convex, normal mosaics, more is known. More specifically, the second part of Lemma 2.4, stating that the upper average number of sides of a cell of such a tiling is at most six is proved in the book [12] of L. Fejes Tóth. The fact that if all cells have at least six sides, then the number of sides with more than six is finite is due to Stehling [24]. Akopyan [2] gave a quantitative estimate for the number of these sides, while Frettlöh et al. [13] showed that the order of magnitude given in [2] is sharp.
We remark that the proofs of the Lemmas 2.1-2.4 (as well as Definition 2.3) can be generalized for all normal mosaics in a straightforward way. Thus, we gave a short proof of the following statement, generalizing the result of Fejes Tóth for any normal mosaic.

### v2

@@ 4. Results about not necessarily polygonal norms
In this section we collect our results about the $\alpha$-honeycomb properties of general norms, and the $\alpha$-Dowker properties of general convex disks. For our first result, recall that $A_{\mathbf{B}^{2}}(n)=n\tan\frac{\pi}{n}$. In the following theorem, we let
Theorem 4.1.
Proof.
$\frac{1}{1+\varepsilon}\left(\frac{n-6}{n-m}\sqrt{m\tan\frac{\pi}{m}}+\frac{6-m}{n-m}\sqrt{n\tan\frac{\pi}{n}}\right)\geq\frac{1}{1-\varepsilon}\sqrt{6\tan\frac{\pi}{6}}.$
Our next result shows that convex disks with sufficiently smooth boundaries ‘asymptotically’ satisfy the log-Dowker property.
@@
$f(n)=\frac{AB(6n^{2}-2)}{(n^{2}-1)^{2}n^{2}}+o\left(\frac{1}{n^{4}}\right),$
@@
In Section 3 we have seen that there are polygons which do not satisfy the $\alpha$-Dowker property for any $\alpha<1$ (see Remark 3.4). Our next theorem can be regarded as a counterpoint of this observation. Before stating it, recall that a convex disk $K$ is smooth if its every boundary point belongs to a unique supporting line of $K$.
@@
First, observe that by our conditions, any supporting line of $K$ intersects $K$ in a unique point, and for any point $p$ of $\operatorname{bd}(K)$ there is a unique supporting line of $K$ that contains $p$. This yields, in particular, that for any $n\geq 3$ and any minimum area convex $n$-gon $Q_{n}$ circumscribed about $K$, $Q_{n}$ has exactly $n$ sides, and the tangent points on those sides are distinct. For any point $p\in\operatorname{bd}(K)$, we denote the unique closed half plane supporting $K$ at $p$ by $H(p)$, and the boundary of $H(p)$ by $L(p)$. We call boundary points of $K$ with opposite unit normal vectors opposite, or antipodal points. Note that if $p,q\in\operatorname{bd}(K)$ are not antipodal points of $K$, there there is a unique closed arc of $\operatorname{bd}(K)$ connecting them that does not contain antipodal points. We denote this arc by $\widehat{pq}$, and denote by $A(\widehat{pq})$ the area of the bounded, connected region $R(\widehat{pq})$ whose boundary consists of $\widehat{pq}$ and the segments of $L(p)$ and $L(q)$ connecting $p,q$ to the intersection point of the two lines. We observe that the problem of finding $A_{K}(n)$ coincides with the problem of finding a tiling of $\operatorname{bd}(K)$ by $n$ arcs $\widehat{x_{i}x_{i+1}}$, $i=1,2,n$, $x_{n+1}=x_{1}$ such that $f=\sum_{i=1}^{n}A(\widehat{x_{i},x_{i+1}})$ is minimal.
First, we fix an arbitrary integer $n\geq 4$, and show that $A_{K}(n-1)+A_{K}(n+1)>2A_{K}(n)$. To do it, we follow the proof of the original theorem of Dowker [8, Theorem 1]
@@
Observe that the family of arcs $\widehat{p_{i}p_{i+1}}$, $\widehat{q_{j}q_{j+1}}$ is a $2$-tiling $\mathcal{T}_{0}$ of $\operatorname{bd}(K)$; that is, every point of $\operatorname{bd}(K)$ belongs to at least two such arcs, and no point belongs to the interior of more than two such arcs. Our main goal will be to modify the $2$-tiling $\mathcal{T}_{0}$ in such a way that the value of $f$ does not increase but the number of covering pairs strictly decreases.
Note that since $\mathcal{T}_{0}$ is the union of two tilings consisting of $(n-1)$ and $(n+1)$ arcs, respectively, $\mathcal{T}_{0}$ contains covering pairs. Assume that $\widehat{v_{i}v_{j}}$ covers $\widehat{v_{k}v_{l}}$. Then let $\mathcal{T}_{1}$ denote the $2$-tiling of $\operatorname{bd}(K)$ in which $\widehat{v_{i}v_{j}}$ and $\widehat{v_{k}v_{l}}$ are replaced by $\widehat{v_{i}v_{l}}$ and $\widehat{v_{k}v_{j}}$. Our main observation, as in the proof of [8, Theorem 1] is that in this case
@@
According to our conditions, $\sum_{S\in\mathcal{T}_{0}}A(S)\geq\sum_{S\in\mathcal{T}_{1}}A(S)$. Furthermore, as $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ are $2$-tilings, and $\widehat{v_{k}v_{l}}$ is already covered twice by the arcs considered in the modification, the number of covering pairs in $\mathcal{T}_{1}$ is strictly less than in $\mathcal{T}_{0}$. Repeating this procedure we obtain a $2$-tiling $\mathcal{T}_{m}$ of $\operatorname{bd}(K)$ for which $\sum_{S\in\mathcal{T}_{0}}A(S)\geq\sum_{S\in\mathcal{T}_{,}}A(S)$ and which does not contain covering pairs. Then, $\mathcal{T}_{m}$ decomposes into two tilings $\{\widehat{v_{1},v_{3}},\widehat{v_{3}v_{5}},\ldots,\widehat{v_{2n-1}v_{1}}\}$ and $\{\widehat{v_{2},v_{4}},\widehat{v_{4}v_{6}},\ldots,\widehat{v_{2n}v_{2}}\}$, each of which contains exactly $n$ arcs.
@@
Since $V_{1}$ and $V_{2}$ are convex $n$-gons circumscribed about $K$, if in (4.2) we have strict inequality, or if $\operatorname{area}(V_{i})>A_{K}(n)$ for $i=1$ or $i=2$, we are done. Thus, in the following we assume for contradiction that $\operatorname{area}(Q_{n-1})+\operatorname{area}(Q_{n+1})=\operatorname{area}(V_{1})+\operatorname{area}(V_{2})$, and $\operatorname{area}(V_{1})=\operatorname{area}(V_{2})=A_{K}(n)$. Then, by the remark right after (4.1) about the equality case in the inequality, there is no arc $\widehat{p_{i}p_{i+1}}$ that contains points $q_{j},q_{j+1}$ in its interior, and vice versa. In particular, there are at least two vertices of $Q_{n-1}$ that coincide with some vertices of $Q_{n+1}$.
@@
Now, let us call a point $q_{j}$ (resp. $p_{i}$) a double vertex, if it coincides with some $p_{i}$ (resp. $q_{j}$). If for some $j$, both $q_{j}$ and $q_{j+1}$ are double vertices, then $q_{j}$ is a double vertex for every $j$ by the previous consideration, which contradicts the fact that $Q_{n+1}$ has strictly more vertices than $Q_{n-1}$. If there is no value of $j$ such that both $q_{j}$ and $q_{j+1}$ are double vertices, then by the previous consideration any arc of $\operatorname{bd}(K)$ connecting two consecutive double vertices contains the same number of vertices from $Q_{n-1}$ and $Q_{n+1}$. But this also contradicts the fact that $Q_{n+1}$ has strictly more vertices than $Q_{n-1}$.
@@ 5. A conjecture of Steinhaus
Furthermore, if $\mathcal{M}$ is a Euclidean plane, then $\mathcal{T}_{hex}$ is a regular hexagonal tiling.
@@
The first part of Theorem 5.2 follows from a straightforward modification of the proof of Theorem 1.9. For the second statement we remark that, according to this proof, the cells of $\mathcal{T}_{hex}$ are homothetic copies of a minimum area hexagon circumscribed about the isoperimetrix of the norm. Thus, if the norm is Euclidean, then these hexagons are regular. ∎
@@
For any normed plane $\mathcal{M}$ there is a hexagonal tiling $\mathcal{T}_{hex}$ of $\mathcal{M}$ such that for any convex tiling $\mathcal{T}$ of $\mathcal{M}$ satisfying $\overline{v}(\mathcal{T})\leq 6$, we have
$\underline{I}(\mathcal{T})\geq I(\mathcal{T}_{hex}).$
Remark 5.4.
We note that Theorem 5.3 remains valid if we replace $\overline{v}(\mathcal{T})$ by $\underline{v}(\mathcal{T})$ and $\underline{I}(\mathcal{T})$ by $\overline{I}(\mathcal{T})$. Furthermore, it is also valid if we replace the condition $\overline{v}(\mathcal{T})\leq 6$ by the property that $\mathcal{T}$ has a cell with at most six sides, and the inequality $\underline{I}(\mathcal{T})\geq I(\mathcal{T}_{hex})$ by $\sup\left\{\frac{\operatorname{perim}_{M}(C)^{2}}{\operatorname{area}(C)}:C\hbox{ is a cell of }\mathcal{T}\right\}\geq I(\mathcal{T}_{hex})$.
@@
Question 5.6.
Let $\mathcal{T}$ be a convex tiling of $\mathbb{R}^{2}$ such that there are universal constants $\lambda,\mu>0$ such that for every cell $C$ of $\mathcal{T}$, we have $\operatorname{ir}(C)\geq\lambda$ and $\frac{\operatorname{cr}(C)}{\operatorname{ir}(C)}\leq\mu$. Is it true that $\overline{v}(\mathcal{T})\leq 6$ (resp. $\underline{v}(\mathcal{T}))\leq 6$? If the answer to both questions is negative, is it true that $\mathcal{T}$ contains a cell with at most six sides?
@@
We define a convex tiling $\mathcal{T}$ of $\mathbb{R}^{2}$ in consecutive steps. More specifically, we start with the family $\mathcal{T}_{0}=\{\mathbf{S}\}$, and for every positive integer $k$, in the $k$th step we add some cells to the family $\mathcal{F}_{k-1}$ of the already defined cells to get the family $\mathcal{F}_{k}$. We will do it in such a way that after each step, $\bigcup\mathcal{F}_{k}=N_{k}\mathbf{S}$ for some positive integer $N_{k}$. Each time we use one of Step A and Step B, described below:
Step A: Let $\bigcup\mathcal{F}_{k-1}=N_{k-1}\mathbf{S}$, and let the sides of this square be $E_{1},E_{2},E_{3},E_{4}$. Consider a rectangle $R$ of side lengths $N_{k-1}$ and $2N_{k-1}$. We set $\mathcal{T}_{k}=\mathcal{T}_{k-1}\cup\{R_{1},R_{2},R_{3},R_{4}\}$, where each $R_{i}$ is congruent to $R$, and one half of a longer side of $R_{i}$ is $E_{i}$. We do it in such a way that the rectangles and $\bigcup\mathcal{T}_{k-1}$ are pairwise nonoverlapping (see Figure 3).
Step B: Let $\bigcup\mathcal{F}_{k-1}=N_{k-1}\mathbf{S}$. We fill the region obtained as the closure of $(N_{k-1}+2)\mathbf{S}\setminus(N_{k-1}\mathbf{S})$ by $44N_{k-1}+$ 4 translates of $\mathbf{S}$.
@@
Fix any positive number $\nu>0$. First, consider the case that we apply Step A $k_{1}$ times, and after that we apply Step B once. Then $N_{k_{1}}=3^{k_{1}}$, and thus, $\mathcal{T}_{k_{1}}$ has $4\cdot 3^{k_{1}}$ boundary edges. In addition, $\mathcal{T}_{k_{1}}$ consists of $14k_{1}+$ 1 cells. We count the number of interior edges: Note that $\mathcal{T}_{0}$ has $4$ edges. Then in each step before the $k_{1}$th one we add $12$ edges, which are all interior edges of $\mathcal{T}_{k_{1}}$. Finally, in the $k_{1}$th step we add four interior edges of $\mathcal{T}_{k_{1}}$. Thus, in total, $\mathcal{T}_{k_{1}}$ has $412k_{1}-$ 4 interior edges. Now, if we set $R_{1}=N_{k_{1}}+\frac{1}{2}$, then
$v_{\square}(R_{1})=\frac{\sum_{C\in\mathcal{T}_{\square}(R_{1})}v(C)}{\operatorname{card}(\mathcal{T}_{\square}(R_{1}))}=\frac{4\cdot 3^{k_{1}}+2(12k_{1}-4)}{4k_{1}+1}.$
@@ 6. An additional remark
Estimating the numbers of sides of cells of tilings has been a longstanding problem of geometry. The classical book of Grünbaum and Shephard [14] shows that it is possible to tile the Euclidean plane with convex heptagons, but that any normal tiling contains infinitely many cells with at most six sides. Here we recall that a (not necessarily convex) tiling $\mathcal{T}$ of $\mathbb{R}^{2}$ is called normal if
@@
    the cells are universally bounded, i.e. there are universal constants $0<\hat{r}<\hat{R}$ such that every cell contains a Euclidean disk of radius $\hat{r}$ and is contained in a Euclidean disk of radius $\hat{R}$.
The statement of Grünbaum and Shephard was reproved in a simpler way by Kazanci and Vince in [16]. For convex, normal mosaics, more is known. More specifically, the second part of Lemma 2.4, stating that the upper average number of sides of a cell of such a tiling is at most six is proved in the book [12] of L. Fejes Tóth in a different way. The fact that if all cells have at least six sides, then the number of cells with more than six sides is finite is due to Stehling [24]. Akopyan [2] gave a quantitative estimate for the number of these sides, while Frettlöh et al. [13] showed that the order of magnitude given in [2] is sharp.
We remark that the proofs of the Lemmas 2.1-2.4 (as well as Definition 2.3) can be generalized for all normal mosaics in a straightforward way. Thus, we gave a short proof of the following statement, generalizing the result of L. Fejes Tóth for any normal mosaic.
@@
7. Acknowledgments
The authors thank Frank Morgan for a valuable observation, stated as Remark 2.8 in the paper.